Find coefficient of correlation, if variance of X series=2.25, standard deviation of Y series=4and regression equation of X on Y is X+0.3Y=1.8
1
Expert's answer
2020-03-11T10:20:05-0400

Formula for coefficient of correlation is the following:

"r=\frac{cov(x,y)}{\sigma_{x}\times \sigma_{x}\times\sigma_{y}\times\sigma_{y}}"

In the regression equation, which is the equation of the type 

"y=a+bx"  , coefficient "b"  has the following meaning:

"b=\frac{cov(x,y)}{\sigma_{x}\times\sigma_{x}}"

Therefore, coefficient of correlation can be written:

"r=\frac{b}{\sigma_{y}\times\sigma_{y}}"

Coefficient b of our regression equation is 

"\frac{1}{0.3}=3.33"

"r=\frac{3.33}{16}=0.208"